1. A random sample of 16 observations is to be drawn from a normal distribution having mean 11 and standard deviation 3.  Let x denote the sample mean.  Find, correct to three decimal places
  2. The probability that x will have a value between 9.2 and 12.2
  3. The value of c for which Px <c=0.03
1
Expert's answer
2021-11-23T17:08:35-0500

(1)

Given that,

mean "=\mu=11"

standard deviation "=\sigma=3"

"n=16 
\\\mu_{\bar{x}}=11 
\\\sigma_{\bar{x}}=\sigma / \sqrt n=3 / 4=0.75"

(2)

"\\ \left.\mathrm{P}(9.2<\bar{x}<12.2)=\mathrm{P}\left[(9.2-11) / 0.75<\left(\bar{x}-\mu_{\bar{x}}\right) / \sigma_{\bar{x}}<(12.2-11) / 0.75\right)\right] 
\\=\mathrm{P}(-2.4<Z<1.6) 
\\=\mathrm{P}(Z<1.6)-\mathrm{P}(Z<-2.4)"

Using z table

=0.9452-0.0082

probability =0.9370

(3)

To find The value of c for which "P(\bar{x}<c)=0.03"

"P(\bar{x}<c)=0.03"

Using standard normal table,

"\begin{aligned}

&P(Z<z)=0.03 \\

&=P(Z<-1.88)=0.03

\end{aligned}"

z = -1.88 Using standard normal z table,

Using z-score formula

"\begin{aligned}

&\bar{x}=\mathrm{z} \times \sigma_{\bar{x}}{+} \mu_{\bar{x}} \\

&\bar{x}=-1.88 \times 0.75+11 \\

&\bar{x}=9.59

\end{aligned}"

So, the value of c =9.59